# 3D Scene Prompting for Scene-Consistent Camera-Controllable Video Generation

### v1

@@ Abstract
We present 3DScenePrompt, a framework that generates the next video chunk from arbitrary-length input while enabling precise camera control and preserving scene consistency. Unlike methods conditioned on a single image or a short clip, we employ dual spatio-temporal conditioning that reformulates context-view referencing across the input video. Our approach conditions on both temporally adjacent frames for motion continuity and spatially adjacent content for scene consistency. However, when generating beyond temporal boundaries, directly using spatially adjacent frames would incorrectly preserve dynamic elements from the past. We address this by introducing a 3D scene memory that represents exclusively the static geometry extracted from the entire input video. To construct this memory, we leverage dynamic SLAM with our newly introduced dynamic masking strategy that explicitly separates static scene geometry from moving elements. The static scene representation can then be projected to any target viewpoint, providing geometrically-consistent warped views that serve as strong 3D spatial prompts while allowing dynamic regions to evolve naturally from temporal context. This enables our model to maintain long-range spatial coherence and precise camera control without sacrificing computational efficiency or motion realism. Extensive experiments demonstrate that our framework significantly outperforms existing methods in scene consistency, camera controllability, and generation quality.
Figure 1: Teaser. Our framework generates the next video chunk that follows a user-specified camera trajectory while maintaining scene consistency. Our dual spatio-temporal conditioning combines the last few frames for temporal continuity and the rendered point cloud for spatial consistency.
@@ 1 Introduction
Camera-controllable video generation (He et al., 2024; Wang et al., 2024b; Jin et al., 2025) aims to synthesize videos following user-specified camera trajectories while maintaining visual coherence and temporal consistency. Recent advances have progressed from generating entirely new videos with controllable viewpoints (Bahmani et al., 2025) to enabling users to extend a single image or short video clips along desired camera paths (He et al., 2024; Agarwal et al., 2025). Yet these methods share a fundamental limitation: they can only process extremely short conditioning sequences, typically just a few frames, which constrains their ability to understand longer videos and hence fails to preserve the rich scene context present in those longer videos. What if we could provide a model with arbitrary-length video sequences and generate continuations that not only follow precise camera controls but also maintain scene consistency with the entire input? Such task, which we refer to as scene-consistent camera-controllable video generation, has immediate applications in film production (Zhang et al., 2025), virtual reality (He et al., 2025), and synthetic data generation (Knapp & Bohacek, 2025).
@@
Scene-consistent, camera-controllable video generation presents three coupled challenges: first, static elements must remain stable across time, while dynamic elements e.g., moving objects or people, should evolve naturally from their most recent states rather than replaying motions from the distant past.Second, effective camera control requires understanding the scene’s 3D geometry: generated content must obey physical constraints, handle occlusions correctly, seamlessly compose dynamic elements with static structure, and plausibly extrapolate previously unobserved regions. Third, these capabilities must remain computationally practical: naïvely conditioning on all input frames scales poorly and becomes intractable as sequence length grows.
@@
Rather than building a bespoke architecture, we leverage strong pretrained video generators, retaining their learned priors and training efficiency, by redesigning how prior content is referenced. Our key insight is to fundamentally rethink this referencing via a dual spatio-temporal conditioning mechanism. Current image-to-video (Yang et al., 2024) and video-to-future-video models  [^1] Throughout our paper, video-to-future-video models refer to models that are capable of generating the subsequent frames of the given input video (e.g., cosmos-predict2 (Agarwal et al., 2025). (Agarwal et al., 2025) achieve realistic generation by conditioning on temporally adjacent frames to maintain short-term consistency and motion continuity. However, adjacency in video is not purely temporal—it can also be spatial. When generating scene-consistent videos, the frames we synthesize may be spatially adjacent to frames from much earlier in the input sequence, particularly when the camera revisits similar viewpoints or explores nearby regions. This dual nature of adjacency suggests a new conditioning paradigm that leverages both temporal and spatial relationships.
@@
Based on these motivations, we propose 3DScenePrompt, a novel video generation framework designed for scene-consistent camera-controllable video synthesis. It takes an arbitrary-length video as context and generates the future video that is consistent with the scene geometry of the context video. The key innovation lies in our dual spatio-temporal conditioning strategy: the model conditions on both temporally adjacent frames (for motion continuity) and spatially adjacent frames (for scene consistency). However, an important consideration for spatial conditioning for our task is that it must provide only the persistent static scene structure while excluding dynamic content, as directly conditioning on spatially adjacent frames from the past would incorrectly preserve dynamic elements. To enable this without temporal contradictions, we construct a 3D scene memory that exclusively represents the static geometry extracted from the entire input video.
@@
To construct this 3D scene memory from dynamic videos, we leverage recent advances in dynamic SLAM frameworks (Zhang et al., 2022; 2024; Li et al., 2024) to estimate camera poses and 3D structure from the input video. To extract only the static regions from the estimated 3D structure, we introduce a dynamic masking strategy that explicitly separates static elements and moving objects. The static-only 3D representation can then be projected to target viewpoints, yielding geometrically-consistent warped views that serve as spatial prompts while allowing dynamic elements to evolve naturally from temporal context alone. Surprisingly, the integration of 3D scene memory provides an additional benefit: the geometrically-consistent warped views provide rich visual references that significantly reduce uncertainty in viewpoint manipulation, enabling precise camera control without any other explicit camera conditioning.
@@
Figure 2: Comparison to existing architectures. (a) Camera-controllable methods condition on a single frame and camera trajectory (Wang et al., 2024b; He et al., 2024; Jin et al., 2025). (b) Video-to-future-video methods use the last few frames of the input video when generating the future video for temporal continuity (Song et al., 2025), but fail to maintain long-term spatial consistency when revisiting viewpoints unseen in the given few frames. (c) Our approach combines temporal conditioning (last few frames) with spatial conditioning (spatially adjacent frames) to achieve scene-consistent generation with precise camera control.
@@ 2 Related Work
Camera-controllable video generation.
Building upon the recent success of video diffusion models (Blattmann et al., 2023; Guo et al., 2023; Yang et al., 2024; Runway, 2024; Brooks et al., 2024), recent works (He et al., 2024; Wang et al., 2024b; Bahmani et al., 2024) have achieved camera-controllable video generation by introducing additional adapters into U-Net-based video diffusion models that accept camera trajectories. For instance, CameraCtrl and VD3D (Bahmani et al., 2024; He et al., 2024) incorporate spatiotemporal camera embeddings, such as Plücker coordinates, via ControlNet-like mechanisms (Zhang et al., 2023). While these methods enable precise trajectory following, they only condition on single starting images, lacking mechanisms to maintain consistency with extended video context. In contrast, our approach enables leveraging entire video sequences as spatial prompts through 3D memory construction, enabling scene-consistent generation that preserves the rich scene context within arbitrary-length inputs.
@@
Recent works (Ren et al., 2025; Yu et al., 2025; Seo et al., 2025) have integrated off-the-shelf geometry estimators into video generation pipelines to improve geometric accuracy. Gen3C (Ren et al., 2025), for instance, similarly adopts dynamic SLAM (Li et al., 2024; Han et al., 2025) to lift videos to 3D representations. However, these methods exclusively address dynamic novel view synthesis—generating new viewpoints within the same temporal window as the input. This constrained setting allows them to simply warp entire scenes without distinguishing static and dynamic elements. Our work fundamentally differs by generating content beyond temporal boundaries, requiring selective masking of dynamic regions during 3D construction—a critical challenge that emerges only when static geometry must persist while dynamics evolve naturally into the future.
@@
Various approaches attempt scene-consistent long video generation through different strategies. ReCamMaster (Bai et al., 2025) and TrajectoryCrafter (Yu et al., 2025) interpolate frames or construct 3D representations but remain confined to the input’s spatiotemporal coverage, essentially performing dynamic novel view synthesis. StarGen (Zhai et al., 2025) scales to long trajectories but assumes static worlds, eliminating temporal dynamics entirely. DFoT (Song et al., 2025) most closely relates to our work, proposing guidance methods that condition on previous frames for scene consistency. However, DFoT also faces fundamental memory constraints when processing extended sequences, limiting its ability to maintain long-range spatial coherence. Our dual spatio-temporal strategy with SLAM-based spatial memory overcomes these limitations by selectively retrieving only the most relevant frames, both temporally and spatially, enabling computationally efficient processing of arbitrary-length videos while maintaining both motion continuity and scene consistency.
@@ 3.1 Problem Formulation and Motivation
where $C_{t}\in\mathbb{SE}(3)$ represents camera extrinsic matrices and $\mathcal{T}$ is a text prompt when a video generator $\mathcal{F}(\cdot)$ is based on pretrained text-to-video priors (Hong et al., 2023; Yang et al., 2024; Bahmani et al., 2025).
@@
This task fundamentally differs from existing video generation paradigms. Existing camera-controllable generation methods (He et al., 2024; Wang et al., 2024b; Bahmani et al., 2024) synthesize videos following user-specified trajectories but only condition on a single image $\mathbf{I}_{\text{ref}}$ or plain text $\mathcal{T}$ (Fig. 2-(a)):
@@
which is insufficient for our task, where the entire underlying 3D scene of the context video should be considered. In contrast, video-to-future-video generation methods such as Cosmos-predict-2 (Agarwal et al., 2025) $\mathcal{G}(\cdot)$ employ temporal sliding windows to generate future frames (Fig. 2-(b)):
@@ 3.2 Towards Scene-Consistent Camera-Controllable Video Generation
The key challenge of scene-consistent camera-controllable video generation lies in reconciling two competing requirements: maintaining consistency with potentially distant frames that share spatial proximity (when the camera returns to similar viewpoints), while evolving dynamic content naturally from the recent temporal context. Ideally, extending the conditioning window of video-to-future-video frameworks to condition on all frames $\mathbf{V}_{\text{in}}$ would ensure optimal global spatial consistency and dynamic smoothness. However, this quickly becomes impractical as the sequence grows, since standard self-attention incurs quadratic time/memory in the sequence length.
@@
Instead of increasing the temporal window size $w$ of the existing video-to-future-video generation methods, we introduce a dual sliding window strategy that conditions on frames selected along both temporal and spatial axes (Fig. 2-(c), and Fig. 3). Beyond the standard temporal window that captures recent motion dynamics, we add a spatial window that retrieves frames sharing similar 3D viewpoints, regardless of their temporal distance:
@@
While this dual conditioning is conceptually appealing, naïvely retrieving and providing spatially adjacent frames directly would be problematic for our task. Since we aim to generate content beyond the input’s temporal boundary, directly conditioning on much earlier frames would freeze outdated dynamics—for example, a walking person seen many frames ago should not reappear at the same location when synthesizing far-future frames. The spatial conditioning should therefore provide only the persistent scene structure while excluding dynamic content. Rather than retrieving individual frames, we introduce a 3D scene memory $\mathcal{M}$ that represents exclusively the static geometry extracted from all spatially relevant frames.
@@
Our 3D scene memory must efficiently encode spatial relationships across all $L$ frames while extracting only persistent static geometry. To construct the 3D scene memory, we leverage dynamic SLAM frameworks (Li et al., 2024; Zhang et al., 2024) to estimate camera poses and reconstruct 3D structure:
@@
where $\hat{\mathbf{C}}=\{\hat{C}_{i}\}_{i=1}^{L}$ are the estimated camera poses, $\mathbf{P}$ represents the aggregated 3D point cloud from the $L$ input frames, and $\mathcal{D}_{\text{SLAM}}(\cdot)$ represents the dynamic SLAM framework. This SLAM integration is effective in that it not only estimates the camera parameters of the input frames but also reconstructs the 3D structure of the scene, which can be further utilized to represent the 3D static geometry in a more efficient manner than other representations (Hong et al., 2024a; b).
@@
We begin with pixel-level motion detection following existing works (Zhang et al., 2024; Han et al., 2025). For each frame pair, we compute optical flow using off-the-shelf models (Wang et al., 2024a; Teed & Deng, 2020; An et al., 2025) ($\text{Flow}_{\text{optical}}$) and compare it against the flow induced by camera motion alone ($\text{Flow}_{\text{warp}}$). Regions where the L1 difference exceeds a specific threshold $\tau$ are marked as potentially dynamic:
$M_{i}^{\text{pixel}}=\mathds{1}\left[\|\text{Flow}_{\text{optical}}-\text{Flow}_{\text{warp}}\|_{1}>\tau\right],$ (6)
where $\mathds{1}[\cdot]$ denotes the element-wise indicator that returns 1 if the condition holds for a pixel and 0 otherwise, i.e., it flags pixels whose optical flow deviates from the warped flow by more than $\tau$.
However, pixel-level detection captures motion only at specific instants and misses complete object boundaries. We therefore propagate these sparse detections to full objects using SAM2 (Ravi et al., 2024), where we sample points from dynamic pixels in the first frame for prompts. Yet this approach still has limitations: static objects that begin moving in later frames may not be detected if they appear static initially.
Our solution employs backward tracking with CoTracker3 (Karaev et al., 2024) to aggregate motion evidence across the entire sequence. From the sampled points in each frame obtained from our pixel-level motion detection, we track these points from all frames back to $t=0$, capturing motions of objects that move at any point. These aggregated points are used to prompt the final SAM2 pass, producing complete object-level masks $M_{i}^{\text{obj}}$ that cleanly separate all dynamic content (Fig. 4-(b)). With the full dynamic mask, we can now obtain the static-only 3D geometry $\mathbf{P}_{\text{static}}$:
@@
Figure 5: Dynamic masking. A three-stage pipeline refines dynamic region detection to produce complete object-level masks: (1) optical-flow differences detect pixel-level motion (Dynamic thresholding); (2) sample points from these regions for all frames and perform backward tracking (BW tracking) with CoTracker3 (Karaev et al., 2024) to aggregate motion evidence across all frames back to t=0t{=}0 (dynamic aggregation), capturing objects that move at any time; (3) propagate aggregated points in the first frame to the entire video using SAM2 (Ravi et al., 2024). The resulting dynamic masks cleanly separate moving elements (people, objects) from the static background, enabling construction of the static-only point cloud $\mathbf{P}_{\text{static}}$.
@@
3.5 Training Dataset Construction
To train our model to effectively handle both static and dynamic scenes while leveraging spatial and temporal conditioning appropriately, we construct training data from diverse video sources with different preprocessing strategies based on scene characteristics.
Data processing.
Our training data comes from two primary sources: RealEstate10K (Zhou et al., 2018) containing static scenes, and OpenVid-1M (Nan et al., 2024), featuring dynamic content. For static videos from RealEstate10K, we directly extract 3D scene geometry without dynamic masking, as the scenes contain negligible motion. In contrast, for dynamic videos from OpenVid-1M, we employ a comprehensive dynamic masking pipeline explained in Section 3.3 to separate static and dynamic elements.
3D scene memory construction.
For effective training, we require videos of sufficient length to construct meaningful 3D scene context. To enable the model to handle arbitrary-length videos as input, we select long videos as the input video with $L>100$. As our video generation framework $\mathcal{F}(\cdot)$ generates $T$ frames, we designate the last $T$ frames as the ground truth future video that the model should generate, while frames before this window serve as input for constructing the 3D scene memory.
Using MegaSAM (Li et al., 2024), we process the entire video to estimate camera poses $\{\hat{\mathbf{C}}_{i}\}_{i=1}^{L}$ for all $L$ frames and reconstruct the 3D structure. We use frames $V_{\text{in}}\in\mathbb{R}^{(L-T)\times H\times W\times 3}$ (i.e., frames before the last $T$ frames) to construct the 3D scene memory, while the camera poses for the last $T$ frames $\{\hat{\mathbf{C}}_{i}\}_{i=L-(T-1)}^{L}$ are used during training to project the scene memory and generate spatial conditioning.
For static scenes, we directly use the reconstructed geometry from $V_{\text{in}}$. For dynamic scenes, we apply our dynamic masks to extract static-only geometry:
$\mathbf{P}_{\text{static}}=\bigcup_{i=1}^{L-T}\mathbf{P}_{i}\odot(1-M^{\text{obj}}_{i})$ (10)
The resulting 3D scene memory $\mathcal{M}=(\hat{\mathbf{C}},\mathbf{P}_{\text{static}})$ provides the spatial context for generating the future $T$ frames. During training, for each target frame $t$ in the last $T$ frames, we use its corresponding camera pose $\hat{\mathbf{C}}_{t}$ to project $\mathbf{P}_{\text{static}}$ from spatially adjacent viewpoints, creating the spatial conditioning Spatial(T)(T) that guides the generation process.
Spatial and temporal conditioning.
To train the model to utilize spatial conditioning effectively, it is crucial to incorporate both static and dynamic regions in our training data. The mixture of RealEstate10K and OpenVid-1M ensures the model tackles scenarios where spatial conditioning should dominate (static regions) and also maintains the prior of generating high-fidelity dynamic motions learned from large-scale pretraining.
For each training sample, we provide:
  • Temporal conditioning: The last $w=9$ frames from the input sequence, capturing recent motion dynamics.
  • Spatial conditioning: Projected views from the static 3D scene memory, synthesized by selecting the point clouds from the top-$n$ spatially adjacent viewpoints based on field-of-view overlap.
This dual conditioning strategy, trained on both static and dynamic content, enables the model to learn when to rely on spatial prompts for scene consistency versus temporal frames for motion continuity. The diversity in our training data—from completely static architectural scenes to videos with dynamic motion—ensures robust generalization to real-world scenarios where both static and dynamic elements coexist.

### v2

@@ Abstract
We present 3DScenePrompt, a framework that generates the next video chunk from arbitrary-length input video chunk while supporting highly complex and precise camera control and preserving scene consistency. Unlike previous methods conditioned on a single image or a short clip, our approach employs dual spatio-temporal conditioning that reformulates context-view referencing across the entire input video. Specifically, we condition on both temporally adjacent frames to ensure motion continuity and spatially adjacent content to preserve scene consistency, enabled by a 3D scene memory that exclusively represents the static geometry extracted from the full input sequence. To construct this memory, we leverage dynamic SLAM with a newly introduced dynamic masking strategy that explicitly separates static scene geometry from moving elements. The resulting static representation can then be projected to arbitrary target viewpoints, providing geometrically consistent warped views that act as strong 3D spatial prompts, while allowing dynamic regions to evolve naturally from temporal context. This design allows our model to maintain long-range spatial coherence and precise camera control without compromising computational efficiency or motion realism. Extensive experiments demonstrate that our framework significantly outperforms existing methods in scene consistency, camera controllability, and generation quality.
Figure 1: Teaser. Our framework generates the next video chunk that follows a user-specified camera trajectory while maintaining scene consistency. Our dual spatio-temporal conditioning jointly leverages the last few frames to ensure temporal continuity and the rendered point cloud to enforce spatial consistency.
@@ 1 Introduction
Camera-controllable video generation (he2024cameractrl; wang2024motionctrl; jin2025flovd) aims to synthesize videos following user-specified camera trajectories while maintaining visual coherence and temporal consistency. Recent advances have progressed from generating entirely new videos with controllable viewpoints (Bahmani_2025_CVPR) to enabling users to extend a single image or short video clips along desired camera paths (he2024cameractrl; agarwal2025cosmos). Yet these methods share a fundamental limitation: they can only process extremely short conditioning sequences, typically just a few frames, which constrains their ability to understand longer videos and hence fails to preserve the rich scene context present in those longer videos. What if we could provide a model with arbitrary-length video sequences and generate continuations that not only follow precise camera controls but also maintain scene consistency with the entire input? Such technology, which we refer to as scene-consistent camera-controllable video generation, has immediate applications in film production (zhang2025generative), virtual reality (he2025pre), and synthetic data generation (knapp2025synthetic).
@@
Scene-consistent camera-controllable video generation poses three intertwined challenges that must be solved jointly. First, static and dynamic elements must be handled differently: while static scene elements should remain consistent throughout generation, dynamic elements such as moving objects and people should evolve naturally from their most recent states rather than rigidly preserving motions from the distant past. Second, camera control demands understanding the underlying 3D geometry of the scene: the generated content must respect physical constraints, properly handle occlusions, and seamlessly compose dynamic elements onto static geometry, while extrapolating plausible content for previously unobserved regions. Third, these capabilities must be achieved within practical computational constraints, as naive approaches that process all input frames quickly become intractable when the input video sequence is long.
@@
How can we tackle this challenging task by leveraging existing video generative models? Our key insight comes from fundamentally rethinking how video models should reference prior content. Current image-to-video (yang2024cogvideox) and video-to-future-video models  [^1] Throughout our paper, video-to-future-video models refer to models that are capable of generating the subsequent frames of the given input video (e.g., cosmos-predict2 (agarwal2025cosmos). (agarwal2025cosmos) achieve realistic generation by conditioning on temporally adjacent frames to maintain short-term consistency and motion continuity. However, adjacency in video is not purely temporal—it can also be spatial. When generating scene-consistent videos, the frames we synthesize may be spatially adjacent to frames from much earlier in the input sequence, particularly when the camera revisits similar viewpoints or explores nearby regions. This dual nature of adjacency suggests a new conditioning paradigm that leverages both temporal and spatial relationships.
@@
Based on these motivations, we propose 3DScenePrompt, a novel video generation framework designed for scene-consistent camera-controllable video synthesis. It takes an arbitrary-length video as context and generates the future video that is consistent with the scene geometry of the context video. The key innovation lies in our dual spatio-temporal conditioning strategy: the model conditions on both temporally adjacent frames (for motion continuity) and spatially adjacent frames (for scene consistency). However, an important consideration for spatial conditioning for our task is that it must provide only the persistent static scene structure while excluding dynamic content, as directly conditioning on spatially adjacent frames from the past would incorrectly preserve dynamic elements. To enable this without temporal contradictions, we construct a 3D scene memory that represents exclusively the static geometry extracted from the entire input video.
@@
To construct this 3D scene memory from dynamic videos, we leverage recent advances in dynamic SLAM frameworks (zhang2022structure; zhang2024monst3r; li2024megasam) to estimate camera poses and 3D structure from the input video. To extract only the static regions from the estimated 3D structure, we introduce a dynamic masking strategy that explicitly separates static elements and moving objects. The static-only 3D representation can then be projected to target viewpoints, yielding geometrically-consistent warped views that serve as spatial prompts while allowing dynamic elements to evolve naturally from temporal context alone. Surprisingly, the integration of 3D scene memory provides an additional benefit: the geometrically-consistent warped views provide rich visual references that significantly reduce uncertainty in viewpoint manipulation, enabling precise camera control without any other explicit camera conditioning.
@@
Figure 2: Comparison to existing architectures. (a) Camera-controllable video generation methods condition on a single frame and camera trajectory (wang2024motionctrl; he2024cameractrl; jin2025flovd). (b) The video-to-future-video generation method (agarwal2025cosmos) leverages the last few frames of the input video to preserve temporal continuity in future video generation; however, it fails to maintain long-term spatial consistency when revisiting previously defined viewpoints that are not visible in the given frames. Unlike these, (c) our approach combines temporal conditioning (last few frames) with spatial conditioning (spatially adjacent frames) to achieve scene-consistent generation with precise camera control.
@@ 2 Related Work
Single-frame conditioned camera-controllable video generation.
Building upon the recent success of video diffusion models (blattmann2023stable; guo2023animatediff; yang2024cogvideox; runway; videoworldsimulators2024), recent works (he2024cameractrl; wang2024motionctrl; bahmani2024vd3d) have achieved camera-controllable video generation by introducing additional adapters into U-Net-based video diffusion models that accept camera trajectories. For instance, CameraCtrl and VD3D (bahmani2024vd3d; he2024cameractrl) incorporate spatiotemporal camera embeddings, such as Plücker coordinates, via ControlNet-like mechanisms (zhang2023adding). While these methods enable precise trajectory following, they only condition on single starting images, lacking mechanisms to maintain consistency with extended video context. In contrast, our approach enables leveraging entire video sequences as spatial prompts through 3D memory construction, enabling scene-consistent generation that preserves the rich scene context within arbitrary-length inputs.
Multi-frame conditioned camera-controllable video generation.
Recently, CameraCtrl2 (he2025cameractrl) and Seaweed-APT2 (lin2025autoregressive) have proposed to take multiple frames as a condition for camera-controllable video generation. This allows the generated videos to maintain temporal smoothness with the provided frames, enhancing the motion fidelity of the generated videos. However, these methods only consider temporal adjacency, which restricts the model from maintaining scene-consistencies with long videos due to memory constraints. In contrast, we introduce SLAM to process the conditioning video to consider dual spatio-temporal adjacency, enabling the model to maintain scene-consistency with long videos under efficient computation.
@@
Recent works (ren2025gen3c; mark2025trajectorycrafter; seo2025vid) have integrated off-the-shelf geometry estimators into video generation pipelines to improve geometric accuracy. Gen3C (ren2025gen3c), for instance, similarly adopts dynamic SLAM to lift videos to 3D representations. However, these methods exclusively address dynamic novel view synthesis—generating new viewpoints within the same temporal window as the input. This constrained setting allows them to simply warp entire scenes without distinguishing static and dynamic elements. Our work fundamentally differs by generating content beyond temporal boundaries, requiring selective masking of dynamic regions during 3D construction—a critical challenge that emerges only when static geometry must persist while dynamics evolve naturally into the future.
@@
Various approaches attempt scene-consistent long video generation through different strategies. ReCamMaster (bai2025recammaster) and TrajectoryCrafter (mark2025trajectorycrafter) interpolate frames or construct 3D representations but remain confined to the input’s spatiotemporal coverage, essentially performing dynamic novel view synthesis. StarGen (zhai2025stargen) scales to long trajectories but assumes static worlds, eliminating temporal dynamics entirely. DFoT (song2025history) most closely relates to our work, proposing guidance methods that condition on previous frames for scene consistency. However, DFoT also faces fundamental memory constraints when processing extended sequences, limiting its ability to maintain long-range spatial coherence. Our dual spatio-temporal strategy with SLAM-based spatial memory overcomes these limitations by selectively retrieving only the most relevant frames, both temporally and spatially, enabling computationally efficient processing of arbitrary-length videos while maintaining both motion continuity and scene consistency.
@@ 3.1 Problem Formulation and Motivation
where $C_{t}\in\mathbb{SE}(3)$ represents camera extrinsic matrices and $\mathcal{T}$ is a text prompt when a video generator $\mathcal{F}(\cdot)$ is based on pretrained text-to-video priors (yang2024cogvideox; Bahmani_2025_CVPR).
@@
This task fundamentally differs from existing video generation paradigms. Existing camera-controllable generation methods (he2024cameractrl; wang2024motionctrl; bahmani2024vd3d) synthesize videos following user-specified trajectories but only condition on a single image $\mathbf{I}_{\text{ref}}$ or plain text $\mathcal{T}$ (Fig. 2-(a)):
@@
which is insufficient for our task, where the entire underlying 3D scene of the context video should be considered. In contrast, video-to-future-video generation methods such as Cosmos-predict-2 (agarwal2025cosmos) $\mathcal{G}(\cdot)$ employ temporal sliding windows to generate future frames (Fig. 2-(b)):
@@ 3.2 Towards Scene-Consistent Camera-Controllable Video Generation
The key challenge of scene-consistent camera-controllable video generation lies in reconciling two competing requirements: maintaining consistency with potentially distant frames that share spatial proximity (when the camera returns to similar viewpoints), while evolving dynamic content naturally from the recent temporal context. Ideally, conditioning on all frames $\mathbf{V}_{\text{in}}$ would ensure optimal global spatial consistency. However, this quickly becomes impractical as the sequence grows, since standard self-attention incurs quadratic time/memory in the sequence length.
@@
Instead of increasing the temporal window size $w$ of the existing video-to-future-video generation methods, we introduce a dual sliding window strategy that conditions on frames selected along both temporal and spatial axes (Fig. 2-(c)). Beyond the standard temporal window that captures recent motion dynamics, we add a spatial window that retrieves frames sharing similar 3D viewpoints, regardless of their temporal distance:
@@
While this dual conditioning is conceptually appealing, naïvely retrieving and providing spatially adjacent frames directly would be problematic for our task. Since we aim to generate future content beyond the input’s temporal boundary, directly conditioning on frames from earlier timestamps would incorrectly preserve dynamic elements (e.g., a walking person from frame 50 should not necessarily reappear at that same location when generating frame 200). The spatial conditioning must therefore provide only the persistent scene structure while excluding dynamic content. Rather than retrieving individual frames, we introduce a 3D scene memory $\mathcal{M}$ that represents exclusively the static geometry extracted from all spatially relevant frames.
@@
Our 3D scene memory must efficiently encode spatial relationships across all $L$ frames while extracting only persistent static geometry. To construct the 3D scene memory, we leverage dynamic SLAM frameworks (li2024megasam; zhang2024monst3r) to estimate camera poses and reconstruct 3D structure:
@@
where $\hat{\mathbf{C}}=\{\hat{C}_{i}\}_{i=1}^{L}$ are the estimated camera poses, $\mathbf{P}$ represents the aggregated 3D point cloud from the $L$ input frames, and $\mathcal{D}_{\text{SLAM}}(\cdot)$ represents the dynamic SLAM framework. This SLAM integration is effective in that it not only estimates the camera parameters of the input frames but also reconstructs the 3D structure of the scene, which can be further utilized to represent the 3D static geometry.
@@
We begin with pixel-level motion detection following MonST3R (zhang2024monst3r). For each frame pair, we compute optical flow using SEA-RAFT (wang2024sea) ($\text{Flow}_{\text{optical}}$) and compare it against the flow induced by camera motion alone ($\text{Flow}_{\text{warp}}$). Regions where the L1 difference exceeds a specific threshold $\tau$ are marked as potentially dynamic:
$M_{i}^{\text{pixel}}=\mathds{1}\left[\|\text{Flow}_{\text{optical}}-\text{Flow}_{\text{warp}}\|_{1}>\tau\right].$ (6)
However, pixel-level detection captures motion only at specific instants and misses complete object boundaries. We therefore propagate these sparse detections to full objects using SAM2 (ravi2024sam), where we sample points from dynamic pixels in the first frame for prompts. Yet this approach still has limitations: static objects that begin moving in later frames may not be detected if they appear static initially.
Our solution employs backward tracking with CoTracker3 (karaev2024cotracker3) to aggregate motion evidence across the entire sequence. From the sampled points in each frame obtained from our pixel-level motion detection, we track these points from all frames back to $t=0$, capturing motions of objects that move at any point. These aggregated points are used to prompt the final SAM2 pass, producing complete object-level masks $M_{i}^{\text{obj}}$ that cleanly separate all dynamic content (Fig. 4-(b)). With the full dynamic mask, we can now obtain the static-only 3D geometry $\mathbf{P}_{\text{static}}$:
@@
Figure 5: Dynamic masking strategy. A three-stage pipeline refines dynamic region detection to produce complete object-level masks: (1) optical-flow differences detect pixel-level motion (Dynamic thresholding); (2) sample points from these regions for all frames and perform backward tracking (BW tracking) with CoTracker3 (karaev2024cotracker3) to aggregate motion evidence across all frames back to t=0t{=}0 (dynamic aggregation), capturing objects that move at any time; (3) propagate aggregated points in the first frame to the entire video using SAM2 (ravi2024sam). The resulting dynamic masks cleanly separate moving elements (people, objects) from the static background, enabling construction of the static-only point cloud $\mathbf{P}_{\text{static}}$.